# Copyright may be transferred without notice, after which this version may no longer be accessible.”

### v1

@@ II SENECA & SENSIM BRIEF
SENECA[24, 21, 22, 23] is a RISC-V-based digital neuromorphic processor designed for SNN/ANN computations in edge applications with limited energy. It handles unstructured spatiotemporal sparsity and irregular data transfers with its optimized neuromorphic coprocessor, three-level memory hierarchy, and dual-controlling system. The processor supports data resolutions of 4b, 8b, and Brain-float16, facilitating advanced learning and optimization. SENECA features an event-based NoC that supports multicasting, source-based routing, and data compression. Its configuration is customizable, including the number of cores, the neural processing elements (NPEs) per core, and optional off-chip memory. SENSIM [19], an open source event-based simulator developed in Python, helps simulate large SNN/ANN models with hardware-aware parameters, providing estimates of energy and latency. It models systems with data transfer between processors in different clock domains and includes parallel processing support for improved execution time. Unlike other simulators such as PyCarl [7], Carlsim [25], and Noxim++ [13], SENSIM offers detailed temporal energy and latency measurements based on hybrid events and clocks and supports a flexible simulation environment. Fig.2 shows an example of SENECA with 64 compute clusters, each comprising an Ibex RISC-V core, memory units, Axon Message Interface (AMI), NoC, Shared Memory Prefetch Unit (SMPU), and NPEs. Fig.2 shows the SENECA Software Development Kit (SDK), which includes the DeltaDNN framework, the SNN simulator, the hardware-aware simulator (SENSIM), the SENECA mapper (SENMap) and the code generator. The hardware details for SENECA are detailed in the prior work [23]. The energy, communication, timing, architecture and simulation parameters extracted from lower-level hardware measurements and used for the experiments are mentioned in the prior work [19].
@@ IV-A Clustering and partitioning
SENMap offers strategies such as partitioning neurons layer-wise, channel-wise, height-wise, and width-wise in a homogeneous and greedy fashion across cores. SENSIM offers clustering layer-wise before distributing them across cores based on the prior information from the analytics framework. In SENSIM, combining layers on a core can lead to interspike distortion and incorrect timestamp updates, leading to erroneous outputs if not handled properly. For example, clustering PilotNet pretrained SNN layers 3 and 4 on the same core can cause deviations, as shown in Figure 5. For smaller SNN models, interspike distortion is minimized by spike counts, but for large scale we suggest measuring accuracy estimated using end signal correlation methods as detailed in section IV-F.
@@ IV-D Bounds and constraints
IV-E Optimization algorithms and problem design
@@
IV-F Inter-spike distortion
@@
IV-G Data collection and visual debugging
@@
IV-H Flexible SNNANN replacement
@@
IV-I Towards closed loop synthesis
@@ V Experimental Results
Parallel meta-heuristics reduced time-to-solution from 3 days to 3 hours. The results indicate that frame rate impacts the precision of PilotNet inference and mapping efficiency, as illustrated in Figures 6,7 & 8. These figures show that increasing frame rates, particularly to 120 fps, affects mapping, making asynchronous transfer ideal for SNNs, where fewer NPEs require additional cores for SIMD arrays. We measure a 40% energy improvement from right to left in Figure 8 as we increase the rate and an optimal mapping requiring just 1 NPE and an increase in cores. In ANNs, rates beyond a threshold can lead to signal interleaving, affecting output accuracy. SENECA’s flexibility as an edge AI chip supports both SNN and ANN needs, with SENMap providing optimal mappings across applications. The correlation values for the SENSIM output signals for the PilotNet dataset as shown in Figure 8 and the values can also be compared here in Table II. The end signals are normalized before correlation and the time shift is estimated in millisecond after scaling up with the time-step in the simulation used.444fps=0 corresponds to a complete event driven system

### v2

@@ II SENECA & SENSIM BRIEF
SENECA[20, 21, 22, 23] is a RISC-V-based digital neuromorphic processor designed for SNN/ANN computations in edge applications with limited energy. It handles unstructured spatiotemporal sparsity and irregular data transfers with its optimized neuromorphic coprocessor, three-level memory hierarchy, and dual-controlling system. The processor supports data resolutions of 4b, 8b, and Brain-float16, facilitating advanced learning and optimization. SENECA features an event-based NoC that supports multicasting, source-based routing, and data compression. Its configuration is customizable, including the number of cores, the neural processing elements (NPEs) per core, and optional off-chip memory. SENSIM [19], an open source event-based simulator developed in Python, helps simulate large SNN/ANN models with hardware-aware parameters, providing estimates of energy and latency. It models systems with data transfer between processors in different clock domains and includes parallel processing support for improved execution time. Unlike other simulators such as PyCarl [7], Carlsim [24], and Noxim++ [13], SENSIM offers detailed temporal energy and latency measurements based on hybrid events and clocks and supports a flexible simulation environment. Fig.2 shows an example of SENECA with 64 compute clusters, each comprising an Ibex RISC-V core, memory units, Axon Message Interface (AMI), NoC, Shared Memory Prefetch Unit (SMPU), and NPEs. Fig.2 shows the SENECA Software Development Kit (SDK), which includes the DeltaDNN framework, the SNN simulator, the hardware-aware simulator (SENSIM), the SENECA mapper (SENMap) and the code generator. The hardware details for SENECA are detailed in the prior work [23]. The energy, communication, timing, architecture, and simulation parameters extracted from lower-level hardware measurements and used for the experiments are mentioned in the prior work [19].
@@ IV-A Clustering and partitioning
SENMap  [^1] Source code of the mapper can be found here. https://github.com/Prithvish04/senMap_paper_submission and https://github.com/Prithvish04/senMap_experiments offers strategies such as partitioning neurons layer-wise, channel-wise, height-wise, and width-wise in a homogeneous and greedy fashion across cores. SENSIM offers clustering layer-wise before distributing them across cores based on the prior information from the analytics framework. In SENSIM, combining layers on a core can lead to interspike distortion and incorrect timestamp updates, leading to erroneous outputs if not handled properly. For example, clustering PilotNet pretrained SNN layers 3 and 4 on the same core can cause deviations, as shown in Figure 5. For smaller SNN models, interspike distortion is minimized by spike counts, but for large scale we suggest measuring accuracy estimated using end signal correlation methods as detailed in section IV-E.
@@ IV-C Bounds and constraints
IV-D Optimization algorithms and problem design
@@
IV-E Inter-spike distortion
@@
IV-F Data collection and visual debugging
@@
IV-G Flexible SNNANN replacement
@@
IV-H Towards closed loop synthesis
@@ V Experimental Results
Parallel meta-heuristics reduced time-to-solution from 3 days to 3 hours. The results indicate that frame rate impacts the precision of PilotNet inference and mapping efficiency, as illustrated in Figures 6,7 & 8. These figures show that increasing frame rates, particularly to 120 fps, affects mapping, making asynchronous transfer ideal for SNNs, where fewer NPEs require additional cores for SIMD arrays. We measure a 40% energy improvement from right to left in Figure 8 as we increase the rate and an optimal mapping requiring just 1 NPE and an increase in cores. In ANNs, rates beyond a threshold can lead to signal interleaving, affecting output accuracy. SENECA’s flexibility as an edge AI chip supports both SNN and ANN needs, with SENMap providing optimal mappings across applications. The correlation values for the SENSIM output signals for the PilotNet dataset as shown in Figure 8 and the values can also be compared here in Table II. The end signals are normalized before correlation and the time shift is estimated in millisecond after scaling up with the time-step in the simulation used.555fps=0 corresponds to a complete event driven system